OpenStudy (anonymous):

a bag is sold for a loss of 10% if it was sold for $45 more, the seller would have gained a profit of 5%. what was the cost price of the bag

OpenStudy (zzr0ck3r):

\(x-0.10x+45=x+0.5x\) solve for \(x\) do you see how I set that up?

OpenStudy (zzr0ck3r):

\(x-0.10x+45=x+0.05x\)

OpenStudy (zzr0ck3r):

typo*

OpenStudy (anonymous):

I think this is correct. \[x=price sold for\] \[y=cost price\] \[x=0.9y \] \[x=1.05y-45\] Substitute first equation value for x into the second equation. Does this work zzzr0ck3r?

OpenStudy (zzr0ck3r):

what you get as an answer?

OpenStudy (anonymous):

300

OpenStudy (zzr0ck3r):

same here

OpenStudy (anonymous):

Nice